cos(-9∏/4)=


cos(-9∏/4)=


Grade:

2 Answers

shivang belwariar
11 years ago

cos(-9∏/4)= cos(9∏/4)=  cos(2∏+∏/4) = cos(∏/4) =cos(45) = 1√2

 

 

PLEASE APPROVE MY ANSWER BY CLICKING YES BELOW THIS .

Shrawan Kher
11 years ago

cos(9pie/4)